Find the strain tensor inside an infinite hollow cylinder of inner and outer radii R₁ and R₂ respectively and deformed by uniform pressure p1 inside and p2 outside.

1
Expert's answer
2022-01-27T15:17:23-0500

"T=\pi \begin{bmatrix}
 p_1 & p_2 \\
 R_1^2 & R_2^2
\end{bmatrix}=\pi(p_1R_2^2-p_2R_1^2)."